### v1

@@ 3.1 Training set
For DBS generation, we subsample structures from databases that are both representative of each material domain and sufficiently extensive, namely MatPES, OC20, OC22, ODAC23, OMol25, and QCML, totaling approximately 125 million structures. From these, we select about 125,000 structures (0.1%) to construct the DBS. For sampling, since training 7net-Omni starts with the weights of 7net-ompa (see the next section), we do not uniformly sample 0.1% from each database. Instead, we assign higher sampling weights to structures that are (i) structurally distant from inorganic systems and (ii) derived from databases whose computational settings differ significantly from the PBE functional used in OMat24 and MPtrj. This is quantified by the force prediction error of 7net-ompa on each database, and the resulting sampling ratios are summarized in Table S1. For ab initio calculations, we employ the Vienna Ab initio Simulation Package (VASP) [104] together with the MPRelaxSet from the pymatgen package to ensure consistency with the computational setups of MPtrj and Alex. We do not apply Hubbard corrections for structures sourced from OC20, ODAC23, OMol25, and QCML.
@@ 3.2 Training
To handle differing atomic reference energies among databases, 7net-MF supports task-wise energy shifts and scaling. Following the approach in ref. [44], we apply only task-wise shifts while maintaining a universal scaling factor across all tasks. During curriculum learning, the shift parameters () for newly added databases are initialized via linear regression using Eq. 8
@@ 3.3 DFT calculation on benchmark datasets
For additional in-house calculations in the below, VASP is used [104] with ENCUT of 520 eV. When required, Grimme’s D3 dispersion correction with Becke-Johnson (BJ) damping is applied [106, 107].
For the noble-metal adsorption benchmark in Fig. 3a, metal slabs with five atomic layers are modeled. Lattice constants are determined from bulk crystal relaxations using an atomic force threshold of 0.01 eV/Å. The pristine slab, adsorbate–slab, and reference molecules are optimized with a force convergence criterion of 0.02 eV/Å. The bottom two layers are fixed during slab and adslab relaxations. A similar procedure is applied to the NEB calculations for the CO2RR process on Pt(111) and Pd(111) surfaces. Specifically, the initial and final states are relaxed with a force threshold of 0.02 eV/Å, followed by a climbing-image NEB calculation using seven intermediate images and a force convergence criterion of 0.05 eV/Å.
@@
For the lattice thermal conductivity benchmark at the r2SCAN fidelity, we follow the computational protocol and ab initio settings described in ref. [96, 97], with two modifications: (i) the XC functional is replaced with r2SCAN, and (ii) the plane-wave cutoff energy is increased from 520 eV to 680 eV, as the r2SCAN functional generally requires a higher cutoff to ensure convergence [108, 109] compared to GGA. To generate displaced structures for second- and third-order force constants and to calculate lattice thermal conductivity, we employ the phono3py package [95].
@@ 4 Data availability
The DBS database and benchmark test results are available on Zenodo at zenodo_doi.
@@ 5 Code availability
The source code for training and trained model parameters of 7net-Omni are available at https://github.com/MDIL-SNU/SevenNet.
@@ 6 Acknowledgements
This work was supported by Neural Processing Research Center program of Samsung Electronics Co., Ltd., and the Virtual Engineering Platform Project of the Ministry of Trade, Industry and Energy (MOTIE) of Korea (grant number: P0022336). The computations were carried out at the Center for Advanced Computations (CAC) at Korea Institute for Advanced Study (KIAS) and Korea Institute of Science and Technology Information (KISTI) National Supercomputing Center (KSC-2025-CRE-0284).
@@ 7 Author contributions
J.K.1, J.Y., and S.H. devised and formalized the idea. J.K.1 and Y.P. developed the training code base. J.K.1, J.Y., Y.L., Y.K.1, J.K.2, H.J., D.H. benchmarked the model performance. S.Y.L., Y.P., and J.W.L. developed and integrated the acceleration of tensor-product operations. J.Y., S.C., and Y.K.2 contributed to the discussion of inference throughput. J.K.1, J.Y., Y.L. and S.H. prepared the manuscript. All authors contributed to discussions and approved the paper. (J.K.1: Jaesun Kim, J.K.2: Jisu Kim, Y.K.1: Yujin Kang, Y.K.2: Yongdeok Kim)

### v2

@@ 3.1 Training set
For DBS generation, we subsample structures from databases that are both representative of each material domain and sufficiently extensive, namely MatPES, OC20, OC22, ODAC23, OMol25, and QCML, totaling approximately 125 million structures. From these, we select about 125,000 structures (0.1%) to construct the DBS. For sampling, since training 7net-Omni starts with the weights of 7net-ompa (see the next section), we do not uniformly sample 0.1% from each database. Instead, we assign higher sampling weights to structures that are (i) structurally distant from inorganic systems and (ii) derived from databases whose computational settings differ significantly from the PBE functional used in OMat24 and MPtrj. This is quantified by the force prediction error of 7net-ompa on each database, and the resulting sampling ratios are summarized in Supplementary Table 1. For ab initio calculations, we employ the Vienna Ab initio Simulation Package (VASP) [108] together with the MPRelaxSet from the pymatgen package to ensure consistency with the computational setups of MPtrj and Alex. We do not apply Hubbard corrections for structures sourced from OC20, ODAC23, OMol25, and QCML.
@@ 3.2 Training
Another key requirement for multi-task training is the alignment of total energies across databases, which differ due to variations in atomic reference energies. Ref. [40] employed atomization energies to align the total energies of the MPtrj and SPICE databases for concurrent mixed-fidelity training of the MACE-Osaka24 model. Similarly, ref. [8] utilized atomization energies and heats of formation to train the multi-fidelity UMA model across heterogeneous datasets. In contrast, ref. [109] adopted a purely statistical approach, fitting atomic reference energies based on elemental composition. This method, applicable even when energies of isolated atom are unavailable for specific ab initio methods, demonstrated improved transferability between GGA and r2SCAN functionals.
Following these studies, the handling of differing atomic energy references across databases was also implemented in 7net-MF, which supports task-wise energy shifts and scaling. Following the observation in ref. [43], we apply only task-wise shifts while maintaining a universal scaling factor across all tasks. In this work, we adopt a modified version of the statistical approach from ref. [109], as calculating isolated-atom energies for all methods represented in Table 1 would be computationally cumbersome. During curriculum learning, the shift parameters () for newly added databases are initialized via linear regression using Eq. 8
@@ 3.3 DFT calculation on benchmark datasets
For additional in-house calculations in the below, VASP [108] is used with ENCUT of 520 eV. When required, Grimme’s D3 dispersion correction with Becke-Johnson (BJ) damping is applied [111, 112].
For the noble-metal adsorption benchmark in Fig. 3a, metal slabs with five atomic layers are modeled. Lattice constants are determined from bulk crystal relaxations using an atomic force threshold of 0.01 eV/Å. The same force convergence criterion is applied for the optimization of pristine slab, adsorbate–slab, and reference molecules. The bottom two layers are fixed during slab and adslab relaxations. A similar procedure is applied to the NEB calculations for the CO2RR process on Pt(111) and Pd(111) surfaces. Specifically, the initial and final states are relaxed with a force threshold of 0.02 eV/Å, followed by a climbing-image NEB calculation using seven intermediate images and a force convergence criterion of 0.05 eV/Å.
@@
For the lattice thermal conductivity benchmark at the r2SCAN fidelity, we follow the computational protocol and ab initio settings described in refs. [98, 99], with two modifications: (i) the XC functional is replaced with r2SCAN, and (ii) the plane-wave cutoff energy is increased from 520 eV to 680 eV, as the r2SCAN functional generally requires a higher cutoff to ensure convergence [113, 114] compared to GGA. To generate displaced structures for second- and third-order force constants and to calculate lattice thermal conductivity, we employ the phono3py package [97].
@@ 4 Data availability
The trained model parameters of 7net-Omni, the DBS database, benchmark test results, and in-house DFT calculation results are available at https://doi.org/10.6084/m9.figshare.30399814.
@@ 5 Code availability
The SevenNet source code for training and evaluation is available at https://github.com/MDIL-SNU/SevenNet.
@@ 6 Acknowledgements
We thank Euitae Lee and Hyuntae Cho for their support in the ML-IAP integration of SevenNet into LAMMPS, and Gijin Kim for discussion on some benchmarks. This work was supported by Neural Processing Research Center program of Samsung Electronics Co., Ltd., and the Virtual Engineering Platform Project of the Ministry of Trade, Industry and Energy (MOTIE) of Korea (grant number: P0022336). The computations were carried out at the Center for Advanced Computations (CAC) at Korea Institute for Advanced Study (KIAS), Korea Institute of Science and Technology Information (KISTI) National Supercomputing Center (KSC-2025-CRE-0284), and computational resources provided by National IT Industry Promotion Agency (RQT-25-070256).
@@ 7 Author contributions
J.K.1, J.Y., and S.H. devised and formalized the idea. J.K.1 and Y.P. developed the training code base. J.K.1, J.Y., Y.L., Y.K.1, J.K.2, H.J., S.J., and D.H. benchmarked the model performance. S.Y.L., Y.P., and J.W.L. developed and integrated the acceleration of tensor-product operations. J.Y., S.C., and Y.K.2 contributed to the discussion of inference throughput. J.K.1, J.Y., Y.L. and S.H. prepared the manuscript. All authors contributed to discussions and approved the paper. (J.K.1: Jaesun Kim, J.K.2: Jisu Kim, Y.K.1: Yujin Kang, Y.K.2: Yongdeok Kim)